### v1

# Yuchen Huang1,  Zhiyuan Fan1,  Zhitao He1,  Sandeep Polisetty2,  Wenyan Li3 Yi R. (May) Fung1 1

Abstract
Pretrained vision-language models (VLMs) such as CLIP excel in multimodal understanding but struggle with contextually relevant fine-grained visual features, making it difficult to distinguish visually similar yet culturally distinct concepts. This limitation stems from the scarcity of high-quality culture-specific datasets, the lack of integrated contextual knowledge, and the absence of hard negatives highlighting subtle distinctions. To address these challenges, we first design a data curation pipeline that leverages open-sourced VLMs and text-to-image diffusion models to construct CulTwin, a synthetic cultural dataset. This dataset consists of paired concept-caption-image triplets, where concepts visually resemble each other but represent different cultural contexts. Then, we fine-tune CLIP on CulTwin to create CultureCLIP, which aligns cultural concepts with contextually enhanced captions and synthetic images through customized contrastive learning, enabling finer cultural differentiation while preserving generalization capabilities. Experiments on culturally relevant benchmarks show that CultureCLIP outperforms the base CLIP, achieving up to a notable 5.49% improvement in fine-grained concept recognition on certain tasks, while preserving CLIP’s original generalization ability, validating the effectiveness of our data synthesis and VLM backbone training paradigm in capturing subtle cultural distinctions. [^1] Our code is publicly available at https://github.com/lukahhcm/CultureCLIP.
1 Introduction
Figure 1: CLIP vs. CultureCLIP. Left: The original CLIP model fails to capture fine-grained contextual visual cues (highlighted by the green box), leading to mismatches with cultural concepts. Although both the image and concept projections fall within the region outlined by the purple dashed box—i.e., the semantic space of Chinese elderly deities—the distance between the image of Yuelao (pink circle) and its correct concept (pink square) is greater than that to an incorrect one such as Taishang Laojun (blue square), i.e., $d_{1}>d_{2}$. Right: CultureCLIP improves fine-grained cultural understanding by jointly aligning concepts with contextualized captions and their corresponding synthetic images, while repelling unrelated concepts and captions in the embedding space (see details in Section 4.2).
Recent advancements in vision-language reasoning (Zhang et al., 2024b; Lu et al., 2024) have revolutionized multimodal understanding by efficiently integrating visual and linguistic semantics within a shared feature space (Jia et al., 2021; Radford et al., 2021). By leveraging large-scale image-text corpora, models such as Contrastive Language-Image Pretraining (CLIP) (Radford et al., 2021) exhibit remarkable generalization capabilities across diverse downstream tasks, including visual question answering (VQA)(Shen et al., 2021; Song et al., 2022), cross-modal retrieval (Koukounas et al., 2024; Baldrati et al., 2023), and zero-shot image classification (Zhou et al., 2022; Saha et al., 2024). CLIP is built on a contrastive learning objective, where two separate encoders are trained to bring matching image-text pairs closer in the feature space while pushing apart non-matching pairs within the same batch. Due to the concise nature of the text in its training data, CLIP is effective at coarse-grained semantic alignment, particularly in identifying the general type of the main object (Radford et al., 2021; Zhang et al., 2024a). However, it often struggles with fine-grained alignment, especially when contextual visual details, such as specific accessories, stylistic cues, or symbolic elements that convey meaning only within particular cultural contexts, are required to distinguish between visually similar but culturally distinct concepts.
For instance, while CLIP may correctly identify both Yuelao (the deity of love and marriage) and Taishang Laojun (a Daoist patriarch) as elderly Chinese deities, it often fails to distinguish between them due to its insensitivity to subtle yet critical visual cues—such as the red thread associated with Yuelao (a symbolic thread that connects lovers), and the alchemy furnace linked to Taishang Laojun (a Taoist alchemy vessel, associated with immortality), as highlighted by the green boxes in Figure 1. These visual details, which carry specific cultural significance, serve as crucial distinguishing cues for humans. A culturally informed observer would instinctively rely on such cues to disambiguate similar-looking yet culturally different figures. This gap leads us to an intriguing question: How can we empower CLIP with contextual cultural knowledge, enabling it to harness fine-grained visual cues for cultural concept differentiation, just like humans do?
A natural approach is to curate a large-scale dataset comprising visually similar cultural concept pairs (i.e., original concepts alongside their hard negatives) accompanied by image-text pairs enriched with contextual cultural knowledge and illustrating subtle visual distinctions. However, collecting such data presents three major challenges: First, high-quality cultural image-text pairs are scarce and costly to annotate. Existing manually curated cultural datasets are typically limited in scale, often containing only a few thousand samples (Nikandrou et al., 2024; Bhatia et al., 2024; Romero et al., 2024), due to the labor-intensive nature of data collection and annotation, as well as the need for domain-specific expertise. While web-scraped data might serve as an alternative (Xu et al., 2023), it frequently introduces substantial noise: images may be ethically inappropriate, copyright-protected, low-resolution, or culturally misleading, and the corresponding text may be mismatched or entirely absent. Second, CLIP’s original training strategy inherently favors concise captions. Specifically, it imposes a 77-token limit, with the majority of alignment achieved within the first 20 tokens (Zhang et al., 2024a). This design makes it particularly challenging to incorporate lengthy, information-rich cultural context directly into the text for training, as doing so may disrupt the original image-text alignment learned by the model. Third, fine-grained hard negatives specifically tailored for visually similar concepts are lacking. In the original CLIP framework, negative samples are randomly drawn within each batch, which significantly reduces the likelihood of including conceptually similar but visually distinct examples. Although recent works (Yuksekgonul et al., 2022; Patel et al., 2024) introduce harder negatives by making minor modifications to captions or images, they still lack structured negative samples that emphasize coarse-grained similarities coupled with fine-grained differences—details that are crucial for enabling culturally grounded visual distinctions.
Motivated by these challenges, we introduce CulTwin, a synthetic dataset of paired concept-caption-image triplets—referred to as Twin Cards — where captions are enriched with cultural background knowledge using a vision-language model, and images are generated by a text-to-image model to reflect fine-grained visual features. Building on CulTwin, we propose CultureCLIP, a contrastive learning framework that jointly aligns concepts, captions, and images in a shared embedding space. As illustrated in Figure 1 (right), our training objective encourages each concept to be attracted to its corresponding caption and image, while repelling it from its culturally contrasting counterpart. Experiments on culturally relevant benchmarks demonstrate that CultureCLIP outperforms the base CLIP model, achieving over a 5% improvement in fine-grained concept recognition for specific tasks. This underscores the effectiveness of our synthetic dataset and training methodology in capturing subtle cultural nuances.
2 Related Work
Advancements in Vision-Language Models
Recent studies on multimodal reasoning (Peng et al., 2025; Zhao et al., 2025; Zhang et al., 2025; He et al., 2025) have significantly advanced vision-language models and broadened their capabilities across various downstream tasks and domains. For instance, Med-Flamingo (Moor et al., 2023) unlocks medical VQA abilities through continued pre-training on paired and interleaved medical image-text data, while ChemVLM (Li et al., 2024), trained on bilingual image-text data, enhances the joint understanding of textual and visual chemical information. In the cultural domain, CultureVLM (Liu et al., 2025) improves cultural understanding by fine-tuning on a large-scale multimodal benchmark, CultureVerse. Despite these advancements, existing vision-language models still struggle to capture fine-grained visual cues and often misclassify visually similar but culturally distinct concepts. We propose CultureCLIP, which aligns cultural concepts with enriched captions and synthetic images through contrastive learning, improving cultural differentiation while preserving generalization capabilities.
Data for Vision-Language Pre-training
Cross-modal mutual information maximization relies on large-scale, diverse training data that captures real-world concepts and relationships. With more than 5 billion internet-derived image-caption pairs, the LAION dataset (Schuhmann et al., 2022) serves as a critical training resource. MetaCLIP (Xu et al., 2023) formalizes CLIP’s implicit data selection via explicit metadata balancing, creating 400M CommonCrawl pairs. SynthCLIP (Hammoud et al., 2024) reduces reliance on web-scraped data by generating over 30 million synthetic pairs. LaCLIP (Fan et al., 2023) enhances text augmentation through in-context language model rewriting. For human-centric AI, fine-grained cultural understanding is essential, yet culturally relevant multimodal data is scarce. While Nayak et al. (2024); Liu et al. (2025) introduced cultural datasets, the diversity of images hinders vision-language models from learning cultural distinctions. In this work, we construct CulTwin, a synthetic dataset comprising concept-caption-image triplets enriched with cultural contextual knowledge.
Contrastive Pre-training
Contrastive learning has become a strong method for multimodal representation learning, with CLIP (Radford et al., 2021) demonstrating scalability and zero-shot transfer potential. More efficient contrastive pre-training methods have been proposed for finer-grained multimodal representations learning (Zhang et al., 2024c; Patel et al., 2024). BLIP-2 (Li et al., 2023) introduces a lightweight Querying Transformer for cost-efficient pre-training. NegCLIP (Yuksekgonul et al., 2022) generates hard negative captions through semantic perturbations. TripletCLIP (Patel et al., 2024) uses hard negative pairs with a triplet contrastive loss. Existing contrastive pre-training methods focus on caption-image pairs and their negative samples, but fail to capture culturally relevant information due to its multidimensional nature. To address this, CultureCLIP enhances cultural understanding by aligning cultural concepts with contextualized captions and synthetic images, while separating unrelated concepts in the embedding space.
3 CulTwin: A Three-stage Cultural Data Curation Pipeline
Figure 2: Left: Data curation pipeline in CulTwin. Right: Architecture of CultureCLIP.
High-quality and diverse data has been shown to be essential for training models like CLIP (Nguyen et al., 2023; Fang et al., 2022). In this section, we present a three-stage data curation pipeline for constructing CulTwin, a synthetic cultural dataset composed of Twin Cards—pairs of concept-caption-image triplets that are visually similar but culturally distinct. The pipeline begins by collecting culturally grounded concepts, their contexts, and visual features, and performing twin matching to identify negative samples that are visually similar but culturally different (Section 3.1). Next, diverse captions are generated by leveraging cultural context and key visual features through a large language model (LLM) (Section 3.2). Finally, images are synthesized from the captions and evaluated using a Vision-Language Model (VLM), which scores each image based on authenticity, consistency, and cultural fidelity to guide data quality filtering (Section 3.3). The resulting concept-caption-image triplets are then organized into Twin Cards for later fine-grained contrastive training. Figure 2 (Left) provides an overview of the full CulTwin data curation pipeline.
3.1 Concept Mining and Twin Matching
We begin with a manually predefined taxonomy covering 229 countries and 8 cultural categories, including Cuisine, Clothing, Animals & Plants, Art, Architecture, Daily Life, Symbols, and Festivals. This taxonomy is designed to capture a broad and representative set of cultural elements (definitions of each category are provided in Appendix A). We then collect culturally meaningful concept candidates through both bottom-up collection and top-down generation.
Bottom-up Concept Collection
In this approach, we first collect candidate concepts and their background information (definitions, images, captions) from Wikipedia. We then use Qwen2.5-VL (Bai et al., 2025) to assess each concept’s cultural relevance, discarding any that are not strongly related to the predefined categories. To ensure strict filtering, we designed prompts (Appendix F) that favor rejection over uncertain inclusion, prioritizing data quality over quantity. For the retained concepts, we assign metadata (country and cultural category) and extract key contextual and visual features.
Top-down Concept Generation
In this approach, we further expand the concept pool using the predefined taxonomy. For each cultural category and country, Qwen2.5-VL is employed to generate culturally grounded concepts, along with associated context and visual features.
After obtaining these filtered concepts from both approaches, we perform twin matching to identify culturally distinct but visually similar hard negatives for each concept, using Qwen2.5-VL conditioned on its context and key visual features. These final concept pairs, together with their metadata and contextual information, form the foundation for generating detailed captions and images in subsequent stages.
3.2 Diverse Caption Generation
In the second stage, we generate diverse, culturally rich captions for each concept in the paired sets. These captions are designed to highlight both key visual elements and cultural nuances, preserving subtle distinctions before image synthesis. We leverage Qwen2.5-VL to incorporate cultural context and salient visual features, producing contextualized descriptions. To mitigate the risk of limited diversity and potential overfitting in synthetic data (Hammoud et al., 2024), we guide the model to vary aspects such as artistic style, scene setting, and compositional details, thereby enriching the visual representation of each concept.
3.3 Image Synthesis and Quality Filtering
Figure 3: Twin Card examples from the Cuisine, Clothing, and Animals & Plants categories. Each Twin Card contrasts two culturally distinct yet visually similar concepts, with each side comprising a concept-caption-image triplet (see Section 3.3). Additional examples are provided in Appendix B.
After caption generation, we synthesize images using Stable Diffusion 3.5 (Rombach et al., 2022), followed by quality filtering with MLLM-as-a-Judge (Chen et al., 2024; Fan et al., 2025a), implemented using Qwen-VL-2.5. Each synthesized image is evaluated across three key dimensions on a scale from 1 to 5:
1) Authenticity: Evaluates the physical realism and adherence to common human understanding. 2) Consistency: Assesses alignment between the image and its caption, ensuring accurate concept representation. 3) Cultural Fidelity: Examines the preservation and correctness of cultural features specific to the concept.
Images receiving a score of 1 in any dimension or an average score below 3 are discarded. To validate the automated filtering, we also perform a human evaluation on a sampled subset, where three PhD-level experts independently score the images using the same criteria. Summary statistics of the automated and human evaluations are shown in Table 1.
Table 1: Filtering Outcomes and Image Quality Scores Across Diverse Cultural Taxonomies
Taxonomy Image Quality Scores After Filtering Filtering Outcomes
MLLM-as-a-Judge Human Evaluation Pass % Retained
Auth $\uparrow$ Cons $\uparrow$ Fid $\uparrow$ Auth $\uparrow$ Cons $\uparrow$ Fid $\uparrow$
Cuisine 4.403 ± 0.528 3.548 ± 0.627 3.750 ± 0.719 4.517 ± 0.866 3.533 ± 0.957 3.550 ± 1.023 77.46 24824/32,046
Clothing 4.469 ± 0.397 3.862 ± 0.422 4.002 ± 0.152 4.550 ± 0.956 3.483 ± 1.218 3.433 ± 1.202 75.52 7250/9,600
Animal & Plants 4.350 ± 0.415 3.979 ± 0.528 4.284 ± 0.366 4.183 ±1.162 3.783 ± 1.185 3.800 ± 1.137 75.63 10966/14,500
Art 4.390 ± 0.444 4.032 ± 0.415 4.093 ± 0.253 4.433 ± 0.761 2.950 ± 1.189 2.917 ± 1.144 72.67 15806/21,750
Architecture 4.222 ± 0.349 3.897 ± 0.402 3.997 ± 0.335 4.183 ± 1.025 3.117 ± 1.112 2.917 ± 1.100 61.32 6377/10,400
Daily Life 4.239 ± 0.436 3.882 ± 0.489 4.152 ± 0.308 4.483 ± 0.904 4.050 ± 1.132 4.100 ± 1.179 72.80 6188/8,500
Symbol 3.802 ± 0.371 3.830 ± 0.291 3.977 ± 0.109 4.467 ± 0.718 3.950 ± 1.310 3.833 ± 1.227 85.12 681/800
Festival 3.971 ± 0.326 3.921 ± 0.332 3.994 ± 0.061 4.167 ± 0.986 3.533 ± 1.087 3.533 ± 1.008 72.12 1731/2,400
Scores are averaged over three dimensions—Auth (image authenticity), Cons (concept consistency), and Fid (cultural fidelity)—rated from 1 to 5. Pass % indicates the proportion of images passing quality thresholds; Retained shows the number of remaining samples per category.
Categories like Cuisine and Clothing exhibit high Authenticity scores, reflecting their tangible and universally recognized nature, making them easier to evaluate accurately by both models and human judges. In contrast, more abstract categories such as Festival tend to have lower Authenticity and Cultural Fidelity scores, particularly in human evaluations, due to their complex and diverse cultural elements. Furthermore, Art and Architecture show a noticeable gap between automated and human evaluations, especially in Consistency and Cultural Fidelity. These categories involve nuanced cultural and conceptual details that automated models struggle to capture, highlighting the importance of human judgment for these intricate domains.
Each final Twin Card is constructed by assembling two triplets—each consisting of a concept, its caption (from Section 3.2), and the corresponding synthesized image—to explicitly highlight cultural contrast while maintaining visual similarity. Example Twin Cards are illustrated in Figure 3, and additional details of CulTwin can be found in Appendix B.
4 CultureCLIP: Fine-Grained Cultural Alignment
4.1 Preliminary
CLIP learns joint image-text representations via an image encoder $\mathcal{F}:\mathcal{I}\rightarrow\mathbb{R}^{d}$ and a text encoder $\mathcal{G}:\mathcal{T}\rightarrow\mathbb{R}^{d}$, projecting inputs into a shared embedding space $\mathcal{V}$ of dimension $d$. Given a batch of $N$ image-text pairs $\{(I_{i},T_{i})\}_{i=1}^{N}$, representations are computed as $\psi_{i}^{I}=\mathcal{F}(I_{i})$ and $\psi_{i}^{T}=\mathcal{G}(T_{i})$. CLIP constructs a similarity matrix $S\in\mathbb{R}^{N\times N}$, where each entry $S_{i,j}=\text{sim}(\psi_{i}^{I},\psi_{j}^{T})$ denotes the cosine similarity between image $I_{i}$ and text $T_{j}$. The contrastive loss encourages each aligned pair (on the diagonal of $S$) to have a higher similarity than all mismatched pairs in the same row or column. Formally:
$\mathcal{L}_{\text{CLIP}}(I,T)=\mathcal{L}_{\text{I2T}}(I,T)+\mathcal{L}_{\text{T2I}}(I,T),$ (1)
$\mathcal{L}_{\text{I2T}}(I,T)=-\frac{1}{N}\sum_{i=1}^{N}\log\frac{\exp(\text{sim}(\psi_{i}^{I},\psi_{i}^{T})/\tau)}{\sum_{k=1}^{N}\exp(\text{sim}(\psi_{i}^{I},\psi_{k}^{T})/\tau)},$ (2)
$\mathcal{L}_{\text{T2I}}(I,T)=-\frac{1}{N}\sum_{i=1}^{N}\log\frac{\exp(\text{sim}(\psi_{i}^{I},\psi_{i}^{T})/\tau)}{\sum_{k=1}^{N}\exp(\text{sim}(\psi_{k}^{I},\psi_{i}^{T})/\tau)}.$ (3)
Here, $\tau$ is a temperature parameter that controls the sharpness of the similarity distribution.
NegCLIP (Yuksekgonul et al., 2022) builds on this framework by introducing hard negative captions $T^{-}$, derived through semantic perturbations of the original texts $T^{+}$. These hard negatives are added alongside the standard in-batch negatives, forming an extended candidate set $\tilde{T}=\{T^{+}\}\cup\{T^{-}\}$ and resulting in a similarity matrix $\tilde{S}\in\mathbb{R}^{N\times 2N}$. The contrastive objective is thus extended beyond standard in-batch negatives to also include explicitly constructed hard negatives:
$\mathcal{L}_{\text{NegCLIP}}(I,T^{+},T^{-})=\mathcal{L}_{\text{I2T\_neg}}(I,T^{+},T^{-})+\mathcal{L}_{\text{T2I}}(I,T^{+}),$ (4)
$\mathcal{L}_{\text{I2T\_neg}}(I,T^{+},T^{-})=-\frac{1}{N}\sum_{i=1}^{N}\log\frac{\exp(\text{sim}(\psi_{i}^{I},\psi_{i}^{T^{+}})/\tau)}{\sum_{k=1}^{N}\exp(\text{sim}(\psi_{i}^{I},\psi_{k}^{T^{+}})/\tau)+\sum_{m=1}^{N}\exp(\text{sim}(\psi_{i}^{I},\psi_{m}^{T^{-}})/\tau)}.$ (5)
TripletCLIP (Patel et al., 2024) further incorporates hard negative images $I^{-}$. The training objective encourages $I^{+}$ to be closer to $T^{+}$ than to $T^{-}$, and symmetrically, $I^{-}$ to align more with $T^{-}$ than with $T^{+}$. This is achieved by summing two NegCLIP-style losses:
$\mathcal{L}_{\text{TripletCLIP}}(I^{+},I^{-},T^{+},T^{-})=\mathcal{L}_{\text{NegCLIP}}(I^{+},T^{+},T^{-})+\mathcal{L}_{\text{NegCLIP}}(I^{-},T^{-},T^{+}).$ (6)
While prior work enhances image-text alignment using modality-specific hard negatives, it tends to focus on coarse-grained semantic differences. We extend this framework by introducing abstract concepts as anchors to better connect detailed captions and images. Through a refined training objective (Section 4.2), our model intends to capture fine-grained cultural semantics with greater precision.
4.2 CultureCLIP
To better capture fine-grained cultural semantics, we build on the Twin Cards introduced in Cultwin, where each card contains two triplets—$(C^{+},T^{+},I^{+})$ and $(C^{-},T^{-},I^{-})$—representing visually similar but culturally distinct concepts. These pairs serve as hard negatives for each other. Our goal is to align each concept with its corresponding caption and image while distinguishing it from its cultural counterpart.
To achieve this, we propose CultureCLIP, a contrastive learning framework that jointly embeds concepts, captions, and images into a shared semantic space. As illustrated in Figure 1, we encourage semantic attraction within each triplet and repulsion across its cultural counterpart. In addition, the overall architecture of our pipeline, including this learning scheme, is shown in Figure 2 (Right). We employ a shared text encoder for both concepts and captions to preserve the original alignment between images and captions while aligning concepts with images. This shared encoder design ensures that fine-tuning for cultural distinctions does not degrade the model’s general cross-modal alignment ability. Our overall training objective is defined as:
$\mathcal{L}_{\text{CultureCLIP}}=\lambda_{c}\cdot\mathcal{L}_{\text{concept}}+\lambda_{t}\cdot\mathcal{L}_{\text{caption}},$ (7)
where $\lambda_{c}$ and $\lambda_{t}$ balance the contributions of concept-level and caption-level objectives. Both objectives are symmetrically formulated to promote attraction within positive triplets and repulsion from their cultural counterparts:
$\mathcal{L}_{\text{caption}}=\mathcal{L}_{\text{NegCLIP}}(I^{+},T^{+},T^{-})+\mathcal{L}_{\text{NegCLIP}}(I^{-},T^{-},T^{+}),$ (8)
$\mathcal{L}_{\text{concept}}=\mathcal{L}_{\text{NegCLIP}}(I^{+},C^{+},C^{-})+\mathcal{L}_{\text{NegCLIP}}(I^{-},C^{-},C^{+}),$ (9)
This structure jointly anchors abstract cultural concepts to specific visual-textual cues, enhancing the model’s ability to differentiate subtle cultural semantics. To further preserve the original model’s generalization ability, we apply the parameter-efficient LoRA method for fine-tuning on this loss, rather than directly training both the visual and text encoders.
5 Experiment
5.1 Experimental Setup
Model Choices and Prompting Strategies
In this work, we leverage Qwen2.5-VL (Bai et al., 2025) as our LLM/VLM (note that the specific choice of model is not the primary focus of this paper). Although more powerful LLMs could potentially further improve data quality, we consider Qwen2.5-VL to strike a good balance between performance and cost, making it a practical choice for our experiments. Given the significant impact that prompt design has on LLM performance (Brown et al., 2020), we meticulously craft prompt templates for each LLM in the pipeline, employing a few-shot learning approach. This includes presenting a set of example input-output pairs to guide the model in various tasks. All prompts used in this study are provided in Appendix F. For the text-to-image generation task, we use Stable Diffusion 3.5 (Rombach et al., 2022) as the model for image synthesis.
Benchmarks
We evaluate the models on both culture-specific and culture-agnostic tasks. To assess cultural understanding, we adapt three benchmarks—GlobalRG-Grounding, GlobalRG-Retrieval (Bhatia et al., 2024), and CROPE (Nikandrou et al., 2024)—into statement-ranking tasks suitable for CLIP-based models. In each task, the model must select the most semantically accurate description for a given image from a set of culturally grounded statements, thereby testing its ability to capture fine-grained, culture-specific visual cues (reported as Accuracy). To assess general vision-language capabilities, we further evaluate the models on MS COCO (Lin et al., 2015) and Flickr30k (Plummer et al., 2016), reporting the average Recall@5 for both image-to-text and text-to-image retrieval tasks. Details of the benchmarks and additional evaluation results on widely used image classification datasets are provided in Appendix D.
Baseline
We first evaluate the performance of CLIP (Radford et al., 2021), NegCLIP (Yuksekgonul et al., 2022), and TripletCLIP (Patel et al., 2024) on our tasks to establish a baseline. Building on this, we introduce CLIP++, NegCLIP++, and TripletCLIP++ as enhanced baselines. In these ++ versions, we train the base CLIP model using our own dataset, aligning the synthetic images with contextualized caption-image pairs, without incorporating the concept.
Implementation Details
All fine-tuned CLIP models use ViT-B/32 (Dosovitskiy et al., 2021) as the image encoder $\mathcal{F}$ and the default CLIP text encoder (Radford et al., 2021) as the text encoder $\mathcal{G}$. We fine-tune the model using a parameter-efficient method, LoRA (Hu et al., 2021), during which $\mathcal{F}$ and $\mathcal{G}$ are frozen, and additional LoRA parameters for these two encoders are applied and trained for 10 epochs with a global batch size of 2048, a learning rate of $3\times 10^{-6}$, weight decay of 0.1, and a cosine learning rate schedule. We employ LoRA primarily to maintain the model’s general capability rather than to reduce memory requirements. In our preliminary experiments with full parameter fine-tuning, we observed significant performance degradation on both culture-specific and culture-agnostic benchmarks, likely due to the distribution gap between the general data used in pretraining and the culture-specific data used during fine-tuning, which disrupted knowledge the model had already acquired. Before inference, LoRA parameters are merged with the backbone transformers, ensuring both efficiency and the preservation of zero-shot transfer capabilities. All models are fine-tuned on 4 Nvidia H20 GPUs using the official Hugging Face Transformers codebase (Wolf et al., 2020).
5.2 Main Results
Evaluation on Culture-Specific Tasks
As summarized in Table 2, CultureCLIP significantly outperforms all baseline models on cultural benchmarks, achieving a 5.49% improvement over CLIP on GlobalRG-G, thus demonstrating strong fine-grained cultural understanding. Directly fine-tuning CLIP on our cultural dataset (CLIP++) leads to a substantial performance drop of 17.93%, highlighting the limitations of naive fine-tuning without explicit negative samples or concept-level alignment. In contrast, our enhanced variants NegCLIP++ and TripletCLIP++ (which incorporate hard negatives) achieve improvements of 2.34% and 0.80% over CLIP, respectively. When further combined with concept-level alignment in CultureCLIP, we observe a large net improvement, emphasizing the effectiveness of our design in leveraging hard negatives and concept supervision for cultural feature discrimination. Similar trends are observed for GlobalRG-R and CROPE. We note that NegCLIP and TripletCLIP are pre-training approaches based on much smaller datasets (e.g., CC3M, CC12M) and do not include culturally relevant data, resulting in substantially lower performance on general benchmarks. For fairness, we do not consider them direct baselines in the cultural evaluation but include them in the table for completeness. All models in Table 2 are trained on the same unfiltered 100k synthetic dataset using LoRA with rank 4 to ensure a fair comparison.
Evaluation on Culture-Agnostic Tasks
On general vision-language tasks, CultureCLIP maintains strong performance and even slightly improves over the baseline, with gains of 0.90% on MS COCO and 0.30% on Flickr30k. This indicates that cultural fine-tuning does not compromise, and may even enhance, general retrieval capabilities.
Table 2: Experimental results on culture-specific and culture-agnostic tasks. All models are trained on the same unfiltered 100k dataset using LoRA with rank 4.
Methods Neg Con Culture-Specific Tasks Culture-Agnostic Tasks
GlobalRG-G GlobalRG-R CROPE MS COCO Flickr30k
CLIP $\times$ $\times$ 63.98 78.22 74.69 65.40 89.0
NegCLIP $\checkmark$ $\times$ - - - 6.50 2.70
TripletCLIP $\checkmark$ $\times$ - - - 10.80 22.00
CLIP++ (ours) $\times$ $\times$ 46.05 49.98 73.62 28.80 50.80
NegCLIP++ (ours) $\checkmark$ $\times$ 66.32 78.41 79.25 65.50 89.20
TripletCLIP++ (ours) $\checkmark$ $\times$ 64.78 78.45 79.25 65.50 89.30
CultureCLIP (ours) $\checkmark$ $\checkmark$ 69.47 78.60 78.84 66.30 89.30
5.3 Ablations
Which contributes more to the model and alignment, caption or concept?
As shown in Table 3, concepts act as abstract semantic anchors, providing stronger cultural alignment capacity and enabling the model to distinguish subtle differences. Captions refine the model’s understanding of specific details but are less critical for recognizing cultural nuances.
Table 3: Ablation study on loss configurations. All models are trained on the same unfiltered 100k dataset using LoRA (rank 4) to preserve general multimodal alignment capabilities.
Configuration $\lambda$ (cap/con) Culture-Specific Tasks Culture-Agnostic Tasks
GlobalRG-G GlobalRG-R CROPE MS COCO Flickr30k
Single Branch (No Negative)
Caption-only w/o neg 1.0 / – 66.95 77.43 79.19 65.60 89.20
Concept-only w/o neg – / 1.0 64.24 77.70 79.19 65.50 89.00
Single Branch (With Negative)
Caption-only w/ neg 1.0 / – 66.27 77.53 79.25 65.50 89.20
Concept-only w/ neg – / 1.0 65.83 77.29 79.19 65.60 88.90
Mixed Branches
Cap (w/o neg) + Con (w/o neg) 0.5 / 0.5 67.29 77.27 79.25 65.50 89.20
Cap (w/ neg) + Con (w/o neg) 0.5 / 0.5 67.12 78.70 79.25 65.60 89.30
Cap (w/o neg) + Con (w/ neg) 0.5 / 0.5 68.81 76.87 79.19 65.50 89.20
Full (Both with Negative)
Both w/ neg (Ours) 0.7 / 0.3 65.93 78.80 79.37 66.80 89.50
Both w/ neg (Ours) 0.5 / 0.5 67.12 78.25 78.60 66.30 89.20
Both w/ neg (Ours) 0.3 / 0.7 69.47 78.60 78.84 66.10 89.30
Can a higher-quality cultural dataset improve performance?
As shown in Table 4, using quality-filtered data (“+QF”), which contains 73.8k high-quality samples after image filtering, consistently improves performance despite using fewer samples compared to the full 100k unfiltered set. For example, comparing Config 5 (+QF, LoRA r=4) to Config 3 (LoRA r=4), and Config 6 (+QF, LoRA r=8) to Config 4 (LoRA r=8), we observe clear performance gains across all cultural benchmarks. This underscores the effectiveness of our data curation pipeline in providing cleaner and more informative supervision for fine-grained cultural alignment.
Table 4: Ablation study on quality filtering (QF) and LoRA. “+QF” uses the 73.8k filtered samples; otherwise, the full 100k dataset is used.
Configuration QF LoRA Rank GlobalRG-G GlobalRG-R CROPE
Baseline $\times$ - 46.95 49.98 73.62
+ QF - 43.64 50.68 74.33
+ LoRA (r=4) $\times$ 4 69.47 78.60 78.84
+ LoRA (r=8) $\times$ 8 65.29 78.03 79.37
+ QF + LoRA (r=4) 4 69.67 78.60 79.21
+ QF + LoRA (r=8) 8 66.48 78.21 79.37
What role does LoRA play in adapting a pretrained model to downstream tasks?
Our ablation results in Table 4 show that without LoRA (e.g., Configs 1 and 2), directly fine-tuning a pretrained model on a domain-specific cultural dataset leads to a substantial performance drop—specifically, 22.52% and 26.03% lower compared to Configs 3 and 5, respectively. This suggests that, without LoRA, the model struggles to effectively absorb specialized supervision, resulting in a significant loss of generalization ability. In contrast, when LoRA is applied (Configs 3–6), the model not only achieves improved performance on cultural benchmarks but also maintains its capabilities on general tasks. These findings highlight LoRA’s essential role in mitigating catastrophic forgetting: it enables the model to flexibly adapt to new cultural signals while preserving the broad vision-language alignment learned during pretraining, ensuring robustness across both domain-specific and general scenarios.
6 Conclusion
In this paper, we introduce CulTwin, a high-quality synthetic dataset of paired concept-caption-image triplets verified by humans, where captions are enriched with cultural background knowledge using a vision-language model, and images are generated by a text-to-image model to reflect fine-grained visual features. Building on CulTwin, we propose CultureCLIP, a novel contrastive learning framework that jointly aligns cultural concepts, captions, and images in a shared embedding space. Our experiments demonstrate that CultureCLIP surpasses baseline models on culture-specific benchmarks, achieving a 5.49% improvement while simultaneously showing performance gains rather than degradation on culture-agnostic benchmarks. These results underscore the effectiveness of our synthetic dataset and training methodology in capturing nuanced cultural distinctions while preserving and even enhancing the model’s generalization capabilities across broader contexts.
7 Limitations and Future Work
While CultureCLIP significantly improves fine-grained cultural understanding, several limitations remain. Both CLIP and CultureCLIP still struggle with cases where the visual distinction is highly abstract or stylistic (see error case analysis in Appendix E). In addition, our current pipeline, for practical considerations, adopts Qwen2.5-VL as an MLLM-as-a-Judge to assess cultural relevance through multidimensional scoring and to guide filtering. However, compared to more advanced models such as GPT-4o, this choice may introduce biases, lead to misjudgments, or lack interpretability. Furthermore, despite the diversity of CulTwin, it is fundamentally a fully synthetic dataset, and there may still exist a distributional gap between synthetic and real-world images. Looking ahead, future work may explore several directions: (1) Improving abstract visual reasoning, by enhancing the model’s capacity to recognize subtle visual cues, such as artistic styles or symbolic meanings; (2) Enhancing the robustness of cultural understanding, by mitigating vulnerability to visual variations and ensuring stable performance across diverse conditions (Fan et al., 2025b); (3) Developing interpretable assessment modules, to enable more robust and transparent cultural data evaluation beyond the current Qwen2.5-VL-based judge; and (4) Bridging the synthetic-real domain gap, by integrating real and synthetic data or applying domain adaptation strategies to improve generalization and visual grounding.
Appendix A Country List and Cultural Taxonomy
In this study, the list of country names is based on data from the GeoNames database: GeoNames.org.
The cultural taxonomy includes the following categories, each representing a significant aspect of cultural identity:
  • Cuisine: Refers to the foods, culinary practices, and cooking methods that are unique to specific regions or cultures. This includes iconic dishes, preparation techniques, and the cultural background behind eating habits, as well as the importance of food in social and religious practices.
  • Clothing: Encompasses traditional garments, accessories, and adornments from various cultures. It includes not only clothing but also items like jewelry, headwear, and footwear that hold cultural significance, reflecting identity, status, and traditions.
  • Animal & Plants: Describes the native species, both fauna and flora, that hold cultural importance. This category includes the use of animals and plants in mythology, cuisine, traditional medicine, and environmental practices, as well as their roles in folklore and symbolism.
  • Art: Includes visual arts, sculptures, and other forms of artistic expression that represent a culture’s aesthetic and artistic heritage. This encompasses paintings, sculptures, performance arts, and crafts that reflect the identity, beliefs, and historical evolution of a community.
  • Architecture: Refers to the design, style, and structures built by a particular culture. This includes traditional houses, temples, monuments, and public buildings that showcase the engineering, material use, and aesthetic values of the culture.
  • Daily Life: Covers the everyday activities, routines, and practices that define how people in a particular culture live. This includes family roles, work habits, and leisure activities, as well as practices around health, education, and community.
  • Symbol: Involves the symbols, logos, and imagery that carry cultural meaning. This category includes national flags, religious icons, mythological figures, and colors that convey beliefs, values, and identity in various contexts.
  • Festival: Encompasses cultural festivals, holidays, and ceremonies, along with the associated customs, rituals, and practices. Examples include events like Chinese New Year, Diwali, and Christmas, each rich in traditions, foods, and rituals that symbolize community and heritage.
Appendix B CulTwin Details
We collected a total of 99,996 Twin Cards, each consisting of two concept–caption–image triplets with culturally distinct negatives. After applying the image quality filtering described in Section 3.3, 73,823 high-quality samples were retained—a scale significantly larger than existing cultural benchmarks such as CROPE (approximately 1k samples). Each concept is represented as a single word, and the corresponding captions have an average length of 14.55 words. All images are synthetically generated from these captions using Stable Diffusion 3.5 Large Turbo, with an efficient generation throughput of approximately 3,000 images per hour on a single H20 GPU. Additional examples of Twin Cards are illustrated in Figure 4.
Figure 4: Additional Twin Card examples showcasing diverse cultural concepts beyond the main text examples, further illustrating fine-grained cultural distinctions and visual similarity.
Appendix C CultureCLIP Pseudocode
def cultureclip_loss(pos_concept_embs, pos_caption_embs, pos_img_embs, neg_concept_embs, neg_caption_embs, neg_img_embs, logit_scale, lambda_caption=0.5, lambda_concept=0.5): """ Full CultureCLIP Loss: Captions negclip + Concepts negclip Aligns concepts with their corresponding captions and images, while distinguishing them from their culturally opposite counterparts. """ # Calculate Captions NegCLIP loss caption_loss_val = ( negclip_loss(pos_img_embs, pos_caption_embs, neg_caption_embs, logit_scale) + negclip_loss(neg_img_embs, neg_caption_embs, pos_caption_embs, logit_scale) ) # Calculate Concepts NegCLIP loss concept_loss_val = ( negclip_loss(pos_img_embs, pos_concept_embs, neg_concept_embs, logit_scale) + negclip_loss(neg_img_embs, neg_concept_embs, pos_concept_embs, logit_scale) ) # Total Loss total_loss = lambda_caption * caption_loss_val + lambda_concept * concept_loss_val return total_loss
Appendix D Benchmark Details and Additional Results
In this section, we provide detailed descriptions of the benchmarks used to evaluate our models, covering both culture-specific and culture-agnostic tasks.
Culture-Specific Tasks
  • GlobalRG-Grounding (Bhatia et al., 2024): Each data point consists of an image, a concept, and a country. We generate four statement-based options for the model to choose from, such as ”The item in the picture is {concept} in {country}.” The correct option corresponds to the appropriate concept for that country, while incorrect options are created by selecting a concept from the same country that does not match the image.
  • GlobalRG-Retrieval (Bhatia et al., 2024): Each data point consists of an image, a category, and a country, without specifying a particular concept. The task focuses on identifying the correct country for the depicted category. Options are phrased as ”The picture depicts a kind of {category} in {country},” with incorrect options generated by randomly substituting the country.
  • CROPE (Nikandrou et al., 2024): The original dataset asks whether the image displays the defined concept (”yes” or ”no”). We filter out ”no” cases where the question concept and definition concept differ, and reformulate the task as a two-choice classification: ”There is {question concept} in the image” or ”There is {definition concept} in the image,” requiring the model to correctly identify the depicted concept.
Culture-Agnostic Tasks
  • MS COCO (Lin et al., 2015): This large-scale dataset comprises over 330,000 images annotated with approximately five captions each. We evaluate using bidirectional retrieval metrics, specifically Text2Image and Image2Text Recall@5, and report their arithmetic mean as the final score.
  • Flickr30k (Plummer et al., 2016): This dataset contains 31,000 images primarily focused on human activities, each paired with five descriptive captions. Similar to MS COCO, we adopt bidirectional retrieval metrics to assess cross-modal alignment.
  • More General Image Classification Benchmarks: To further assess generalization capabilities, we evaluate our models on several widely used image classification benchmarks, including FER2013, ImageNet-1k, ImageNet-A, ImageNet-O, ImageNet-R, VOC2007, CIFAR-10, and CIFAR-100. These datasets collectively cover a broad spectrum of visual tasks, such as facial expression recognition, large-scale object classification, out-of-distribution robustness, multi-label classification, and both coarse- and fine-grained category recognition. Empirical results demonstrate that CultureCLIP consistently maintains, and in some cases slightly improves, performance on these benchmarks. This suggests that the incorporation of cultural fine-tuning does not compromise general vision-language alignment or classification capabilities but rather enhances overall robustness and versatility. A comprehensive summary of these results, including Top-1 Accuracy (Acc1), Top-5 Accuracy (Acc5), and Mean Per-Class Recall (MPCR), is provided in Table 5.
Table 5: Performance on general image classification benchmarks (%). We report Top-1 Accuracy (Acc1), Top-5 Accuracy (Acc5), and Mean Per-Class Recall (MPCR) across various datasets. CultureCLIP maintains or slightly improves general performance despite additional cultural training.
Model FER2013 ImageNet-1k ImageNet-A ImageNet-O ImageNet-R VOC2007 CIFAR-10 CIFAR-100
Top-1 Accuracy (Acc1)
OpenAI CLIP 41.22 63.37 31.51 47.55 69.31 76.45 89.77 64.24
Caption (w/o neg) 41.57 63.37 31.61 47.80 69.28 76.52 89.84 64.48
Caption (w/o neg) + Concept (w/o neg) 41.20 63.35 31.59 47.65 69.33 76.42 89.86 64.40
Caption (w/ neg) 41.59 63.39 31.61 47.65 69.26 76.52 89.80 64.45
CultureCLIP 41.26 63.37 31.60 47.65 69.31 76.48 89.83 64.47
Top-5 Accuracy (Acc5)
OpenAI CLIP 94.78 88.82 64.23 78.30 88.81 95.93 99.61 88.78
Caption (w/o neg) 94.68 88.82 64.31 78.10 88.80 95.98 99.62 88.91
Caption (w/o neg) + Concept (w/o neg) 94.82 88.83 64.20 78.25 88.00 95.94 99.63 88.84
Caption (w/ neg) 94.69 88.82 64.33 78.10 88.79 95.96 99.63 88.87
CultureCLIP 94.80 88.80 64.29 78.10 88.85 95.95 99.63 88.85
Mean Per-Class Recall (MPCR)
OpenAI CLIP 36.10 63.36 32.63 48.85 67.92 80.59 89.83 64.21
Caption (w/o neg) 36.40 63.36 32.75 49.09 67.92 80.60 89.86 64.44
Caption (w/o neg) + Concept (w/o neg) 36.00 63.37 32.72 49.01 67.94 80.64 89.87 64.40
Caption (w/ neg) 36.42 63.39 32.72 49.04 67.87 80.58 89.83 64.41
CultureCLIP 36.21 63.37 32.77 49.00 67.96 80.60 89.85 64.41
Appendix E Error Case Analysis
Distinguishing between Gongbi and Xieyi styles
One illustrative failure case involves differentiating between two classic Chinese painting styles: gongbi and xieyi. The gongbi style is known for its meticulous brushwork, fine lines, and realistic details, often used to depict flowers, birds, and other subjects in a highly controlled and precise manner. In contrast, the xieyi style (literally ”writing ideas”) emphasizes freehand expression, bold strokes, and abstract or suggestive forms rather than realistic details. In this example, both CLIP and CultureCLIP misclassified an input gongbi painting as xieyi. However, CultureCLIP assigned a lower confidence to the incorrect label (72% vs. CLIP’s 78%), indicating a modest calibration improvement. This suggests that while our model still struggles with highly abstract stylistic distinctions, it demonstrates better uncertainty awareness compared to the original CLIP, which is a step toward more nuanced cultural reasoning.
Figure 5: Example failure case comparing CLIP and CultureCLIP on a gongbi painting. While both models misclassified it as xieyi, CultureCLIP exhibited lower confidence in its wrong prediction, suggesting better calibration.
Appendix F Prompt Engineering
Bottom-up Filtering Prompt for Qwen2.5-VL Please determine whether the concept ’{title}’ clearly and unambiguously belongs to one of the eight cultural categories (Cuisine, Clothing, Animals & Plants, Art, Architecture, Daily Life, Symbol, or Festival). If the concept is only loosely related, culturally ambiguous, or does not strongly align with any of the categories, please select A to ensure strict filtering. The concept does not clearly belong to any of the eight categories. The concept is clearly and directly related to one of the eight categories. Answer only with A or B’. Output the result in JSON format as follows: { "concept_type": "A" or "B", }
Figure 6: Bottom-up filtering prompt for Qwen2.5-VL.
Bottom-up Classification Prompt for Qwen2.5-VL Given a cultural concept, definition, definition caption, and definition image, your task is to classify and extract the following information in English: - Country: The country or region associated with the concept. - Cultural Category: IMPORTANT - You MUST choose exactly ONE category from these eight predefined categories: * Cuisine * Clothing * Animal & Plants * Art * Architecture * Daily Life * Symbol * Festival Do not use any other categories. If the concept doesnt clearly fit into one of these categories, choose the closest match. - Context: A short 20-word description in English that provides insight into the cultural and functional use of the concept. - Key Visual Features: The visual features that distinguish the concept (e.g., shape, material, color, size) described in English. Examples: Input: Concept: Kimono Definition: A traditional Japanese garment with long, wide sleeves and an obi sash, worn for formal occasions and festivals. Definition Caption: A colorful kimono with crane patterns, displayed at Kyotos textile museum. Definition Image: example_image.jpg Output: Country: Japan Category: Clothing Context: A traditional Japanese garment with long, wide sleeves and an obi sash, worn for formal occasions and festivals. Key Visual Features: Long, wide sleeves, obi sash, colorful patterns, traditional Japanese style. Current Task: Concept: {concept} Definition: {definition} Definition Caption: {caption} Definition Image: {image_url} Generate the Country, Cultural Category (MUST be one of the eight predefined categories), Context, and Key Visual Features in English:
Figure 7: Bottom-up classification prompt for Qwen2.5-VL.
Top-down Generation Prompt for Qwen2.5-VL Given a country and cultural category, your task is to generate the following: - Concept: A cultural concept that fits the provided country and category. - Context: A short 20-word description that provides insight into the cultural and functional use of the concept. - Key Visual Features: The visual features that distinguish the concept (e.g., shape, material, color, size). Examples: Input: Country: China Cultural Category: Food Output: Concept: Mantou Context: Steamed wheat bun symbolizing prosperity and wisdom, commonly served during festivals and family gatherings. Key Visual Features: Pillowy white appearance, round or rectangular shape, smooth surface, typically palm-sized. Input: Country: Japan Cultural Category: Art Output: Concept: Ukiyo-e Context: Traditional woodblock prints depicting scenes from everyday life, nature, and historical events. Key Visual Features: Flat color blocks, bold outlines, vibrant pigments, rectangular format, detailed patterns. Current Task: Country: {country} Cultural Category: {category} Generate the concept, context, and key visual features:
Figure 8: Top-down generation prompt for Qwen2.5-VL.
Twin Matching Prompt for Qwen2.5-VL Given a cultural concept from a certain category, its context, and its visual features, your task is to generate the following: - New Concept: A visually similar but culturally different concept from the same category. - New Context: A short 20-word description in English that provides insight into the cultural and functional use of the generated concept. - New Key Visual Features: The visual features that distinguish the concept (e.g., shape, material, color, size) from the original concept. Examples: Input: Category: Art Concept: Erhu Context: A two-stringed Chinese musical instrument played with a bow, often used in traditional Chinese music. Key Visual Features: Two strings, a bow, a wooden body, and a horsehair bow. Output: New Concept: Guzheng New Context: A Chinese zither-like instrument with a large, rectangular wooden body and multiple strings, played with plucking. New Key Visual Features: A large, rectangular wooden body, multiple strings, and a plucking mechanism. Current Task: Category: {category} Concept: {concept} Context: {context} Visual Features: {visual_features} Generate a new concept, context, and visual features:
Figure 9: Twin matching prompt for Qwen2.5-VL.
Diverse Caption Generation Prompt for Qwen2.5-VL Given a cultural concept, context, and key visual features, your task is to generate 10 different captions that describe the concept in various scenarios while preserving its cultural significance. The captions must reflect different styles or settings, but they should all clearly include the key visual features. Follow these guidelines: - Emphasize the key visual differences in the scenes (e.g., shape, size, cooking method, setting). - Retain the cultural or functional context (e.g., everyday use, ceremonial purpose, tradition). - Ensure the differentiating features (e.g., shape, texture, size, material, use) are clearly reflected in each caption. - Each caption should be under 15 words. - Each caption should be unique, showing different perspectives or settings, but should always include the key visual features. Examples: Input: Concept: Xiaolongbao (Soup Dumplings) Context: Steamed soup dumplings with delicate, thin wrappers, filled with savory broth and pork, typically steamed in bamboo baskets. Key Visual Features: Delicate, thin wrappers and steamed in bamboo baskets. Output: 1. A chef carefully places Xiaolongbao in a bamboo steamer, showcasing their thin, translucent wrappers and savory broth inside. 2. A steaming bamboo basket of Xiaolongbao, delicate wrappers holding savory broth, served in an elegant Shanghai restaurant. 3. Steamed xiaolongbao resting in bamboo baskets, ready to be served during a family meal. 4. Crispy fried xiaolongbao, golden-brown and served with dipping sauce, sitting in bamboo baskets. 5. Miniature xiaolongbao filled with crab roe, elegantly presented in bamboo baskets at a Cantonese restaurant. 6. Steaming xiaolongbao with delicate skin, served in bamboo baskets during a traditional Chinese New Year meal. 7. Bamboo-steamed xiaolongbao, filled with savory broth, served alongside hot tea in a Beijing teahouse. 8. Translucent, plump xiaolongbao, freshly steamed in bamboo baskets for a cozy brunch setting. 9. Steamed xiaolongbao with pork filling, served in bamboo baskets with chili oil at a street food stall. 10. Elegant xiaolongbao, arranged in bamboo baskets, presented at a lavish festive feast. Current Task: Concept: {concept} Context: {context} Key Visual Features: {visual_features} Generate 10 different captions, each reflecting a different style or scene, but all incorporating the key visual features:
Figure 10: Diverse caption generation prompt for Qwen2.5-VL.
Prompt for Quality Evaluation on Authenticity Please analyze this image and rate its authenticity on a scale of 1 to 5. You can refer to the context to help you make the decision. Focus on whether the concept shown is realistic and follows common sense. Consider: - Are all elements anatomically and physically correct? - Does everything look natural and possible in the real world? - Are there any unrealistic or deformed features? Examples: Input: Concept: Erhu Context: The Erhu is a traditional Chinese musical instrument played with a bow. It consists of a wooden body and two strings, and is known for its expressive and resonant sound. Often used in Chinese classical and folk music, it is typically performed in both solo and ensemble settings. Image: An image showing a person playing Erhu with three hands in mid air Output: 1 Input: Concept: Erhu Context: The Erhu is a traditional Chinese musical instrument played with a bow. It consists of a wooden body and two strings, and is known for its expressive and resonant sound. Often used in Chinese classical and folk music, it is typically performed in both solo and ensemble settings. Image: An image showing a person playing Erhu with two hands sitting on the chair; Output: 5 Now, give you the image, concept and its corresponding context: Concept: {concept} Context: {context} Output only the score:
Figure 11: Prompt for quality evaluation on authenticity.
Prompt for Quality Evaluation on Consistency Please analyze this image and rate its consistency with the concept on a scale of 1 to 5. You can refer to the context to help you make the decision. Focus on whether the image accurately depicts the specified concept without showing wrong concepts. Consider: - Does the image show exactly the concept mentioned? - Are there any mismatched or wrong elements? - Is the concept clearly and accurately represented? Examples: Input: Concept: Tang Sancai Context: A type of Chinese glazed pottery from the Tang Dynasty, known for its colored glazes (green, yellow, and brown) often used for tomb figurines and decorative pieces. Image: An image showing a blue-and-white porcelain bowl Output: 1 Input: Concept: Tang Sancai Context: A type of Chinese glazed pottery from the Tang Dynasty, known for its colored glazes (green, yellow, and brown) often used for tomb figurines and decorative pieces. Image: An image showing exactly a Tang Sancai horse Output: 5 Now, give you the image, concept and its corresponding context: Concept: {concept} Context: {context} Output only the score:
Figure 12: Prompt for quality evaluation on consistency.
Prompt for Quality Evaluation on Cultural Fidelity Please analyze this image and rate its cultural fidelity on a scale of 1 to 5. Focus on whether the cultural elements are accurate and appropriate for the specific context. Consider: - Are all cultural elements accurate for this context? - Are there any mixed or incorrect cultural elements? - Does everything align with the cultural background specified? Examples: Input: Concept: Mexican Day of the Dead Context: A Mexican holiday celebrating deceased loved ones, featuring marigold flowers, sugar skulls, and candle altars. Image: An image showing marigold flowers, skull paintings, and candle altars. Output: 5 Input: Concept: Mexican Day of the Dead Context: A Mexican holiday celebrating deceased loved ones, featuring marigold flowers, sugar skulls, and candle altars. Image: An image showing showing marigold flowers, skull paintings, and Chinese joss paper money. Output: 1 Now, give you the image, concept and its corresponding context: Concept: {concept} Context: {context} Image:{image} Output only the score:
Figure 13: Prompt for quality evaluation on cultural fidelity.